HOW MUCH SALT MUST BE ADDED TO 60 KG OF A 20% SOLUTION OF SALT TO INCREASEE IT TO A 40% SOLUTION OF SALD
Answer by stanbon(57967)

HOW MUCH SALT MUST BE ADDED TO 60 KG OF A 20% SOLUTION OF SALT TO INCREASEE IT TO A 40% SOLUTION OF SALD
-----
Equation:
salt + salt = salt
0.20*60 + 1*x = 0.40(60+x)
-----
20*60 + 100x = 40*60 + 40x
60x = 20*60
x = 20 KG (amt of salt that must be added)
==================
Cheers,
Stan H.